# MAROON: A Framework for the Joint Characterization of Near-Field High-Resolution Radar and Optical Depth Imaging Techniques

### v2

@@ 6.2.1. Depth Deviation at Single and Multiple Object Distances
Additionally, we illustrate the distribution of the depth deviation across all objects for varying placement distances of 30 cm, 40 cm, and 50 cm. The results are displayed as boxplots in Figure 5.
Metric Type Cardboard Sponge Scrubber Plushie Tape Dispenser Statue
RF ToF $\text{C}_{\text{g}}$ 0.13 ($\pm$ 0.06) 1.52 ($\pm$ 0.97) 0.58 ($\pm$ 0.29) 0.81 ($\pm$ 0.47) 0.31 ($\pm$ 0.23) 0.27 ($\pm$ 0.25)
@@
Metric Type S1 Hand Open Hand Printed Flat Mirror Candle Flowerpot (Transparent) V1 Metal Plate
RF ToF $\text{C}_{\text{g}}$ 0.36 ($\pm$ 0.38) 0.71 ($\pm$ 0.78) 0.87 ($\pm$ 0.26) 1.50 ($\pm$ 1.12) 1.31 ($\pm$ 1.21) 0.12 ($\pm$ 0.05)
@@
Table 4. We measure the sensor-GT deviation with respect to $\text{C}_{\text{g}}$, $\text{C}_{\text{s}}$, P, and $\text{P}_{\text{e}}$, which we list in the form $(\mu\pm\sigma)$, consisting of the mean $\mu$ and standard deviation $\sigma$ in centimeters, computed over the entire metric domain, respectively. The best results among all sensors of one metric type are highlighted in bold and the worst results are underlined. The results are discussed in Section 7.
Figure 5. We plot the distribution of the mean error across all objects with respect to different object-to-sensor distances. Solid (—) and dashed (- -) horizontal lines indicate the median and the mean of the distribution, respectively. The results are discussed in Section 7.
@@
Figure 6. We present the received signal magnitude (mean absolute phasor value) in relation to the mean depth deviation of metric $\text{C}_{\text{g}}$ and further investigate both quantities with respect to varying object material, geometry (median surface incidence angle), and size (relative surface area). We highlight measurements where large objects appear outside the radar’s field of view (FOV) in grey regions, as they exhibit comparatively higher depth deviations compared to the ground truth reconstructions that possibly extend beyond this FOV. The results are discussed in Section 9.
6.2.2. RF Signal Response at Single Object Distance
@@
In addition to the phase shift, the point-wise confidence of the backprojection algorithm for RF ToF is significantly affected by the signal amplitude; however, disentangling amplitude and phase in the presence of signal interference imposes the same challenges that arise from recovering the phase shift itself. To avoid inducing additional bias through the assumptions made in signal processing, we investigate the unprocessed signal response in the MIMO imaging radar across multiple objects. For each object at 30 cm object-to-sensor distance, we compute the mean absolute value out of all complex phasors received from the raw signal of size $\text{N}_{\text{RX}}\times\text{N}_{\text{TX}}\times N_{f}$, averaged over 10 frames. We refer to this quantity as signal magnitude, emphasizing the difference from signal amplitude.
@@
To investigate correlations between signal response and reconstruction quality, we analyzed the signal magnitude relative to the depth deviation from ground truth, expressed as the mean of $\text{C}_{\text{g}}$. These findings are shown on the left in Figure 6. We further differentiate between signal magnitude (upper row) and depth deviation (bottom row), examining their relationships with respect to object material, geometry, and size from left to right. Object materials are categorized into six classes, with detailed information available in the supplementary material. The goal of this classification is to highlight material differences on a coarse level, noting the large object variety that still persists within one material class. The object geometry is quantified by the median angle in degrees between the point-wise surface normals of the GT reconstruction and the depth direction (along the $z$-axis) of the imaging radar. As we positioned the objects to align their primary orientation with the viewing direction of the planar square-shaped antenna aperture — which particularly becomes important for flat objects — the median angle mainly reflects geometric complexity, with objects having a higher surface incidence angle showing larger portions oriented away. Object size is determined by the relative surface area compared to the radar antenna aperture. It is computed from the fraction of the object’s 2D axis-aligned bounding box $A$ (in the $x$- and $y$-axis) inside of the 2D axis-aligned bounding box $B$ of the antenna array by using the formula: $(A\cap B)/B$.
7. Sensor Characterization by Depth Deviation
First, we present complementary perspectives on the data, where we put the depth deviation of all 45 objects into relation with the different metric types and, subsequently, the different types of depth imagers. We analyze each representation in turn, highlighting common trends, in combination with previously stated results of Section 6.2. Finally, we examine the depth deviation obtained specifically for the selected objects of Table 4.
7.1. General Trends
Interpreting the extensive numerical data on depth deviations in Table 4 and the supplementary material can be challenging, so we provide visual, complementary views in this section. In order to relate different quantities to each other, we use barycentric interpolation based on triples of metric types (Figure 7) and sensors (Figure 8), respectively. For each triple $(\mu_{a},\mu_{b},\mu_{c})$, the mean values for depth deviation ($\mu$) of each object are transformed to affine coordinates $(w_{a},w_{b},w_{c})$ by using the formula $w_{\{a,b,c\}}=\mu_{\{a,b,c\}}/(\mu_{a}+\mu_{b}+\mu_{c})$. Circle locations closer to a triangle corner indicate higher relative depth deviation. Moreover, as the triples are drawn from a set of four, the triangles are arranged in the shape of an unfolded tetrahedron, highlighting that each triangle’s contents can be seen as a projection of barycentric coordinates within a (3D) tetrahedron $w_{\{a,b,c,d\}}=w_{\{a,b,c,d\}}/(w_{a}+\cdots+w_{d})$.
Figure 7. A complementary view on the depth deviation across different metric types. For each triplet of metrics $M_{i}\in\{\text{$\text{C}_{\text{g}}$},\text{$\text{C}_{\text{s}}$},\text{P},\text{$\text{P}_{\text{e}}$}\}$, we convert each mean depth deviation $\mu_{i}$ to affine coordinates, $\bar{\mu}_{i}=\mu_{i}/\sum_{i}\mu_{i}$, that map an object’s errors in to a triangle whose corners correspond to metrics $M_{i}$. All 45 MAROON objects are shown as circles, with selected objects from Table 4 highlighted in solid colors. Samples closer to a triangle corner indicate a higher relative depth deviation in the corresponding metric.
Figure 8. A complementary view on the depth deviation across different sensors. The sensors are denoted as R (RF ToF), N (NIR ToF), A (Active Stereo), and P (Passive Stereo), respectively. Analogously to Figure 7, we convert the mean depth deviations $\mu_{i}$ to affine coordinates within triangles corresponding to all possible sensor triples. All 45 MAROON objects are shown as circles, with selected objects from Table 4 highlighted in solid colors. Samples closer to a triangle corner indicate a higher relative depth deviation for the corresponding sensor.
In Figure 7, we provide a qualitative comparison of each sensor’s depth deviation from GT with respect to the four presented metrics. In dense reconstructions, as is typical for optical depth sensors, metrics based on nearest neigbors (here, $\text{C}_{\text{g}}$ and $\text{C}_{\text{s}}$) are bound to be lower than those based on projection (P and $\text{P}_{\text{e}}$); they also tend to be more resilient against noise. For RF reconstructions, however, that are prone to sparse depth maps, Chamfer distances often create false matches; accordingly, $\text{C}_{\text{g}}$ dominates for RF ToF compared to the other metrics. For further discussion regarding the sparsity of RF reconstructions, see Section 9.
For optical sensors, a marginal trend towards the corners of $\text{C}_{\text{s}}$ and P, away from the silhouette-resilient $\text{P}_{\text{e}}$ and $\text{C}_{\text{g}}$, cf. Table 3, indicates the presence of noise at object silhouettes.
In Figure 8, we observe a considerable spread of depth deviations across different sensors. As noted in Section 6.2, the relative depth deviation between sensors ranges from 1.9 to 3.4 mm, that is, the variation in the depicted normalized error occurs within a comparatively small range of absolute errors. As a general trend, the two stereo sensors have the lowest depth deviation (see triangles N–R–A and N–R–W), with a moderate edge for active stereo, particularly for metrics $P$ and $P_{\text{e}}$, see the $A$$P$ axis, where passive stereo‘s performance degrades. This is consistent with the results in Section 6.2.1, where we find that the active stereo has the highest number of best results. However, given that relative depth deviations between sensors differ by only a few millimeters, we conclude that the seemingly improved depth quality of active stereo sensors is of minimal significance. Moreover, an examination of the scatter plot reveals a marginal weight trend towards the corners of the two ToF sensors. Reconstruction errors between the two ToF sensors are highly object-dependent, and many objects, including many of the highlighted ones, demonstrate multi-path effects due to perfect signal reflections, retro-reflectivity, or partial signal transmission. Further details on this will be discussed in the next section.
Considering the results over varying object-to-sensor distances in Figure 5, we observe that the depth deviation of the passive stereo and the NIR ToF sensor is significantly more distance-specific compared to active stereo and RF ToF. For passive stereo, this may be due to a decrease in effective spatial resolution, where $\delta_{z}$ directly depends on $\delta_{x,y}$ (see supplementary). As the distance between the object and the sensor increases, $\delta_{x,y}$ decreases, resulting in a loss of high-frequency color details while the object appears smaller in the image. Compared to passive stereo, the active stereo sensor has a comparably higher effective resolution, assuming that the resolution of both sensors differs in accordance with $\delta_{xyz}$ in Table 1. The unique active NIR pattern may also be less sensitive to decreases in spatial resolution, maintaining the quality of correspondence matches. The trend for the NIR ToF sensor aligns with findings from Bamji et al. (2018) for 30–50 cm distances. However, we argue that absolute errors for a target with 20% reflectivity do not fully represent all objects in our experiments. We suggest that, in addition to the expected decrease in spatial resolution, a greater depth deviation with increasing distance arises from the signal-to-noise ratio with respect to environmental light, which typically decreases over distance due to the inverse-square law.
Notably, in our results, RF ToF does not exhibit a distance-dependent depth deviation, unlike its optical counterpart. This seems to contradict Table 1, where $\delta_{z}$ for RF ToF degrades more rapidly with depth, compared to optical sensors; however, the theoretical decay with depth stems from the worsening separability of neighboring point targets, which is not a pertinent scenario in our database, as the recorded targets primarily have smooth surfaces at locations where valid reconstructions are measured. Moreover, our setup minimizes mmWave interactions with external objects, limiting noise primarily to the object itself. As a result, the signal-to-noise ratio of RF ToF is considerably less sensitive to changes in object-to-sensor distance compared to NIR ToF, assuming no interference from external sources.
@@ 7.1. General Trends
Figure 9. For selected objects, we show the reconstructed pointclouds (left) next to their deviation from to the MVS reconstruction (right). The signed depth deviation P* is given for each pixel $(u,v)$ in centimeters. All measurements in the domain $\bm{M}^{+}(u,v)$ are projected onto the GT reconstruction and mapped to color using a combination of a symmetrical logarithmic scale and linear mapping between $[-0.5,0.5]$. Furthermore, we quantify the mean deviation of P* in centimeters below each sensor measurement.
7.2. Object-specific Observations
@@
The following section will analyze the objects in Table 4 in regard to their relative depth deviations over one or multiple sensors. For RF ToF, we find that the least deviation relative to the mean of all metrics occurs with planar object geometries (V1 Metal Plate, Cardboard), followed by more complex shapes (Statue, S1 Hand Open, Hand Printed Flat). For a deeper discussion, see Section 9.
@@
For similar reasons, NIR ToF shows large depth deviations for visually transparent objects like Flowerpot, Candle, Sponge, and Tape Dispenser. Although both RF ToF and NIR ToF are susceptible to multi-path effects, our experiments suggest these effects do not occur for the same objects. Further examination of wavelength-specific multi-path effects, with a particular focus on partially transmissive materials, will be discussed in Section 8.
Additional sources of high depth deviation for NIR ToF include thin structures (Scrubber), which reduce the sensor’s effective spatial resolution. Highly reflective objects (Metal Plate) may cause sensor oversaturation, while perfectly specular materials (Mirror) yield depth values not from the object’s surface but from the first weak scatterer encountered along the reflection path.
For the active stereo sensor, we observe higher depth deviations for textureless and partially transmissive materials (Sponge, Candle). Similar to NIR ToF, the uniqueness of the active NIR light pattern can be compromised by multi-path effects. Finally, as expected, the passive stereo camera is particularly sensitive to textureless objects (Sponge, Candle).
8. Discussion ToF in Partially Transmissive Media
@@
In the remainder, we will now study four selected objects, shown in Figure 9. In addition to Table 4, this figure visualizes depth reconstruction deviations using a signed version P* of our metric P, color-encoded on a symmetrical logarithmic scale (SymLogNorm [^2] https://matplotlib.org/3.8.4/api/_as_gen/matplotlib.colors.SymLogNorm.html), with a linear mapping between $[-0.5,0.5]$ centimeters. The supplementary material includes signed versions of P and $\text{P}_{\text{e}}$ for all MAROON objects.
In the NIR domain, the most prominent effect of global transport occurs for objects with strong internal scattering. Here, the ToF reconstructions exhibit systematic depth deviations of P*, generally biased toward larger distances than the ground truth. This is consistent with the light traveling an additional distance due to scattering within the object before being remitted again, so that the observed propagation time of the actively transmitted signal is consistently longer than for a direct (local) reflection at the object surface. Examples in Figure 9 for internal scattering include subsurface scattering (S1 Hand Open, Sponge, Candle) and inter-reflections within hollow objects (Tape Dispenser). Extended path length due to subsurface scattering is an established effect, systematically measured by Lukinsone et al. (2020). For human skin (as in our dataset S1 Hand Open), and for points of incidence and exitance one millimeter apart, Lukinsone et al. observe effective sub-surface path lengths of up to $26\pm 3$ mm at 800 nm wavelength, which — in the context of a ToF sensor — would result in a systematic depth deviation of half that path length ($\approx+13$ mm). At the same time, however, for human skin a significant portion of the total remitted light stays very close to the point of incidence (Jensen et al., 2001), suggesting that the bulk of the received signal experiences even smaller path length extensions, lending plausibility to our measured systematic depth deviation of $+4.9$ mm for S1 Hand Open to be due to subsurface scattering.
@@
9. Discussion of MIMO Radar Signal and Reconstruction
@@
Our observations in Section 7 suggest that RF ToF reconstructions are generally less complete than those of optical sensors, as illustrated in the second row of Figure 9, where quality is best when surface orientations align with the antenna aperture. To identify sources of incomplete RF ToF reconstructions, we first analyze the raw signal response — without inducing additional bias from the reconstruction algorithms — and subsequently relate it to the quality of the measured depth after reconstruction.
@@
In Figure 6, we presented the received signal magnitude across objects with varying material, geometry (surface incidence angle) and size (relative surface area). Following the scatter plot order from left to right, we will now discuss common trends, noting that it remains challenging to disentangle the presented quantities, as the large variability across objects prevents us from isolating one quantity while keeping the others constant. Further results where we discuss the potential relation between object size and object geometry are given in the supplementary material.
@@
Considering object material, metal and metallic-coated objects generally show higher signal magnitudes, which is consistent with previous studies (Ahmed, 2021, 2014). With a considerably lower spread, large magnitudes are also observed for captures of human skin, which is highly reflective due to its rich water content (Ahmed, 2014). Object materials made of polymers, fibers and stone, or foam generally respond with much smaller signal magnitude.
@@
Regarding object geometry and size, no significant global trends are observed; however, consistent patterns emerge within subsets of the same material, particularly in the metal and polymer classes, which have the highest number of samples. For objects of more complex geometry, with a median surface incidence angle greater than $10^{\circ}$, large portions of their surface area face away from the antenna array, resulting in decreased signal responses compared to planar objects aligned closely with the antenna aperture (¡$10^{\circ}$). The reflection direction of the transmitted signal depends on the surface normal’s orientation. As the angle between this normal and the depth axis increases, the solid angle of the object relative to the planar square-shaped antenna aperture (cf. Table 1) decreases. In other words, a decreasing area around the hemisphere of outgoing reflection directions is aligned with the $65^{\circ}$ field of view of the RX antennas, resulting in reduced signal energy reception, and thus radar cross-section (Knott et al., 2004). Superficially this resembles the well-known cosine law in radiometry, but the exact quantitative relationship depends on the object’s location relative to the individual RX and TX antennas and is further modulated by the non-trivial radiation and signal lobes of the antennas.
@@
Aside from object geometry, the received signal magnitude also appears to increase with object size for non-metal materials. Within these material classes, the highest signal magnitude is achieved for objects close to or even larger (objects partially outside FOV) than the antenna aperture. As the latter also typically exhibits a low median surface incidence angle, it remains questionable, whether this observation can be attributed to object size or object geometry. To address this, we additionally visualize the relation between the two quantities in the supplementary material. Assuming that signal magnitude is proportional to the received energy, our findings correspond to the fact that, the reflected signal energy received at the RX antennas directly depends on the surface area of the irradiated object, in case the energy density is constant. To summarize, we identified object geometry and material as two factors that influence the maximum signal magnitude at each receiver, while also observing indications of the relevance of object size.
We now continue to investigate the relationship between signal magnitude and the observed depth deviation, which is illustrated in the upper left scatter plot in Figure 6. Focusing on polymer and metal objects, we generally find no direct relationship between signal magnitude and depth deviation. Polymer and metal objects have a large spread in the $x$- and $y$-axis, respectively, while the opposite axis has comparably low variation. The received signal magnitude may have a more significant influence on the reconstruction quality in less constrained scenarios, involving multiple objects and signal sources, where, for example, depth filtering becomes increasingly relevant. In our experiments, within the RF near field, we suggest that reconstruction quality is more closely related to the distribution of received signals across antennas, influenced primarily by the local antenna layout and characteristics.

### v4

@@ 6.2.1. Depth Deviation
Additionally, we illustrate the distribution of the depth deviation across all objects for varying placement distances of 30 cm, 40 cm, and 50 cm in the supplementary material.
Metric
Type
Cardboard Sponge Scrubber Plushie Tape Dispenser Statue
RF ToF
$\text{C}_{\text{g}}$
0.13 ($\pm$ 0.06) 1.52 ($\pm$ 0.97) 0.58 ($\pm$ 0.29) 0.81 ($\pm$ 0.47) 0.31 ($\pm$ 0.23) 0.27 ($\pm$ 0.25)
@@
Metric
Type
S1 Hand Open Hand Printed Flat Mirror Candle Flowerpot (Transparent) V1 Metal Plate
RF ToF
$\text{C}_{\text{g}}$
0.36 ($\pm$ 0.38) 0.71 ($\pm$ 0.78) 0.87 ($\pm$ 0.26) 1.50 ($\pm$ 1.12) 1.31 ($\pm$ 1.21) 0.12 ($\pm$ 0.05)
@@
Table 4. We measure the depth deviation with respect to $\text{C}_{\text{g}}$, $\text{C}_{\text{s}}$, P, and $\text{P}_{\text{e}}$, which we list in the form $(\mu\pm\sigma)$, consisting of the mean $\mu$ and standard deviation $\sigma$ in centimeters, computed over the entire metric domain, respectively. The best results among all sensors of one metric type are highlighted in bold and the worst results are underlined. The results are discussed in Section 15.
@@
Figure 5. On the left, object material, geometry (median surface incidence angle), and size (relative surface area) are put in relation to received signal response (mean signal magnitude, top row) and mean depth deviation (bottom row). On the right, both quantities are directly compared to each other. Measurements, where large objects appear outside the radar’s antenna aperture, are highlighted in gray regions, as they exhibit higher depth deviations compared to the ground-truth reconstructions, which may extend beyond this aperture; this is attributed to the comparably small field of view and the surface reflection characteristics with respect to radio waves (see supp. mat.). The results are discussed in Section 7.3.
6.2.2. RF Signal Response
@@
The point-wise confidence of the backprojection algorithm for RF ToF is considerably affected by the signal amplitude; however, disentangling amplitude from phase in the presence of signal interference imposes the same challenges that arise from recovering the phase shift itself. To avoid inducing additional bias through the assumptions made in signal processing, we investigate the unprocessed signal response of the MIMO imaging radar across multiple objects. For each object at 30 cm object-to-sensor distance, we compute the mean absolute value out of all complex phasors received from the raw signal, averaged over 10 frames. We refer to this quantity as signal (phasor) magnitude, emphasizing the difference from signal amplitude.
Differentiating between signal response and reconstruction quality, we examine their relationships with respect to object material, geometry, and size. Our findings are shown on the left of Figure 5, where we visualize these relationships for signal magnitude (upper row) and mean depth deviation (bottom row) in isolation. On the right, we further investigate correlations between signal magnitude and mean depth deviation.
Object materials are categorized into six classes, with detailed information available in the supplementary material. The goal of this classification is to highlight material differences on a coarse level, noting the large object variety that still persists within one material class.
@@
The object geometry is quantified by the median angle in degrees between the point-wise surface normals of the GT reconstruction and the depth direction (along the $z$-axis) of the imaging radar. As we positioned the objects to align their primary orientation with the viewing direction of the planar square-shaped antenna aperture — which particularly becomes important for flat objects — the median angle mainly reflects geometric complexity, with objects having a higher surface incidence angle showing larger portions oriented away. An extension of Figure 5 (left) is available in the supplementary material, visualizing the correlation of the median angle as well as per-angle measurements with respect to the depth deviation of all four sensors.
Object size is determined by the relative surface area compared to the radar antenna aperture. It is computed from the fraction of the object’s 2D axis-aligned bounding box $A$ (in the $x$- and $y$-axis) inside of the 2D axis-aligned bounding box $B$ of the antenna array by using the formula: $(A\cap B)/B$.
@@
Figure 6. For selected objects, we show the reconstructed point clouds (left) next to their deviation from to the MVS reconstruction (right). The signed depth deviation P* is given for each pixel $(u,v)$ in centimeters. All measurements in the domain $\bm{M}^{+}(u,v)$ are projected onto the GT reconstruction and mapped to color using a combination of a symmetrical logarithmic scale and linear mapping between [0.5,0.5][-0.5,0.5] centimeters. The mean deviation of P* is quantified in centimeters below each sensor measurement.
7. Discussion
In this section, we provide a general discussion of the previously reported results related to depth deviation, followed by two focused discussions of time-of-flight sensor effects that offer complementary perspectives on these results. Regarding the latter, we investigate effects of partially transmissive media and explore RF ToF as a particularly under-explored sensor technology, focusing on the received signal response in relation to depth deviation.
We note that a comprehensive sensor characterization, highlighting common trends across all 45 objects, is provided in the supplementary material, where we discuss the interpretation of metrics, the depth deviation over varying distances, as well as relative depth deviations between sensors.
7.1. Discussion of Object-specific Depth Deviation
The following section will analyze the objects in Table 4 in regard to their relative depth deviations over one or multiple sensors.
Radio-Frequency Time-of-Flight
@@ 7.1. Discussion of Object-specific Depth Deviation
For RF ToF, we find that the least deviation relative to the mean of all metrics occurs with planar object geometries (V1 Metal Plate, Cardboard), followed by more complex shapes (Statue, S1 Hand Open, Hand Printed Flat). For a deeper discussion, see Section 7.3.
@@
Near-infrared Time-of-Flight
For similar reasons, NIR ToF shows large depth deviations for visually transparent objects like Flowerpot, Candle, Sponge, and Tape Dispenser. Both RF ToF and NIR ToF are susceptible to multi-path effects; however, our experiments suggest these effects do not occur for the same objects. Further examination of wavelength-specific multi-path effects, with a particular focus on partially transmissive materials, will be discussed in Section 7.2.
Additional sources of high depth deviation for NIR ToF include thin structures (Scrubber), which reduce the sensor’s effective spatial resolution. Highly reflective objects (Metal Plate) may cause sensor oversaturation, while perfectly specular materials (Mirror) yield depth values from the first weak scatterer after perfect reflection.
Active and Passive Stereo Sensors
For the active stereo sensor, we observe higher depth deviations for textureless and partially transmissive materials (Sponge, Candle). Similar to NIR ToF, the uniqueness of the active NIR light pattern can be compromised by multi-path effects. The passive stereo camera is particularly sensitive to textureless objects (Sponge, Candle).
7.2. Discussion of Time-of-Flight Sensors: Partially Transmissive Media
@@
In the remainder, we will now study the four selected objects in Figure 6. In addition to Table 4, this figure visualizes depth deviations using a signed version P* of metric P, color-encoded on a symmetrical logarithmic scale (SymLogNorm [^1] https://matplotlib.org/3.8.4/api/_as_gen/matplotlib.colors.SymLogNorm.html), with a linear mapping between [0.5,0.5][-0.5,0.5] centimeters. The supplementary material includes signed versions of P and $\text{P}_{\text{e}}$ for all MAROON objects.
Near-infrared Time-of-Flight
In the NIR domain, the most prominent effect of global transport occurs for objects with strong internal scattering. Here, the ToF reconstructions exhibit systematic depth deviations of P*, generally biased toward larger distances than the ground truth. This is consistent with the light traveling an additional distance due to scattering within the object before being remitted again, so that the observed propagation time of the actively transmitted signal is consistently longer than for a direct (local) reflection at the object surface. Examples in Figure 6 for internal scattering include subsurface scattering (S1 Hand Open, Sponge, Candle) and inter-reflections within hollow objects (Tape Dispenser). Extended path length due to subsurface scattering is an established effect, systematically measured by Lukinsone et al. (2020). For human skin (e.g. S1 Hand Open), and for points of incidence and exitance one millimeter apart, Lukinsone et al. observe effective sub-surface path lengths of up to $26\pm 3$ mm at 800 nm wavelength, which — in the context of a ToF sensor — would result in a systematic depth deviation of half that path length ($\approx+13$ mm). At the same time, however, for human skin a significant portion of the total remitted light stays very close to the point of incidence (Jensen et al., 2001), suggesting that the bulk of the received signal experiences even smaller path length extensions, lending plausibility to our measured systematic depth deviation of +4.9+4.9 mm for S1 Hand Open to be due to subsurface scattering.
Radio-Frequency Time-of-Flight
@@
7.3. Discussion of MIMO Radar: Signal Response and Depth Deviation
@@
Our observations in Section 15 suggest that RF ToF reconstructions are generally less complete than those of optical sensors, as illustrated in the second row of Figure 6, where depth deviations are lowest when surface orientations align with the antenna aperture.
Initially, this seems to contradict the expectation that larger antenna apertures should capture more surface compared to cameras, given the variety of positions and viewing angles from the individual RX-TX antenna pairs; this advantage, however, seems to be mitigated by the fact that most object surface reflections appear to be specular (Lu et al., 2013). This means that reflections at surfaces oblique to the aperture are only received by a small fraction of antennas, thereby weakly contributing to the overall signal response, potentially at the same level as noise.
We discuss further sources of incomplete RF ToF reconstructions in the next sections, where we first analyze the raw signal response — without inducing additional bias from the reconstruction algorithms — and subsequently relate it to the quality of the measured depth after reconstruction.
7.3.1. Radio-Frequency Signal Response
@@
In Figure 5 (top left), we presented the received signal magnitude across objects with varying material, geometry and size. Following the scatter plot order from left to right, we will now discuss common trends, noting that it remains challenging to disentangle the presented quantities, as the large variability across objects prevents us from isolating one quantity while keeping the others constant.
Influence of Material
@@
Metal and metallic-coated objects generally show higher signal magnitudes, which is consistent with previous studies (Ahmed, 2021, 2014). With a considerably lower spread, large magnitudes are also observed for captures of human skin, which is highly reflective due to its rich water content (Ahmed, 2014). Object materials made of polymers, fibers and stone, or foam generally respond with much smaller signal magnitude.
Regarding object geometry and size, no significant global trends are observed; however, consistent patterns emerge within subsets of the same material, particularly in the metal and polymer classes, which have the highest number of samples. We will discuss these patterns within the next paragraphs.
Influence of Geometry
@@
For objects of more complex geometry, with a median surface incidence angle greater than $10^{\circ}$, large portions of their surface area face away from the antenna array, resulting in decreased signal responses compared to planar objects aligned closely with the antenna aperture ($10^{\circ}$). The reflection direction of the transmitted signal depends on the surface normal’s orientation. As the angle between this normal and the depth axis increases, the solid angle of the object relative to the planar square-shaped antenna aperture (cf. Table 1) decreases. In other words, a decreasing area around the hemisphere of outgoing reflection directions is aligned with the approximate, $53^{\circ}$ field of view of the RX antennas, resulting in reduced signal energy reception, and thus radar cross-section (Knott et al., 2004). Superficially this resembles the well-known cosine law in radiometry, but the exact quantitative relationship depends on the object’s location relative to the individual RX and TX antennas and is further modulated by the non-trivial radiation and signal lobes of the antennas.
Influence of Size
@@
Aside from object geometry, the received signal magnitude also appears to increase with object size for non-metal materials. Within these material classes, the highest signal magnitude is achieved for objects close to or even larger than the antenna aperture. As the latter also typically exhibits a low median surface incidence angle, it remains questionable, whether this observation can be attributed to object size or object geometry. To address this, we additionally visualize the relation between the two quantities in the supplementary material. Assuming that signal magnitude is proportional to the received energy, our findings correspond to the fact that, the reflected signal energy received at the RX antennas directly depends on the surface area of the irradiated object, in case the energy density is constant.
7.3.2. Radio-Frequency Depth Deviation
Following the previous section, we now summarize the results of Figure 5 (bottom left), where we relate the depth deviation to varying object material, size, and geometry in their respective scatter plots.
Influence of Material
Similar to our findings for signal response, metal objects generally exhibit the lowest depth deviation with a relatively small spread compared to other material classes, indicating that object material influences reconstruction quality.
Influence of Size and Geometry